The current value of a firm is 425,100 dollars and it is 100% equity financed. The firm is considering restructuring so that it is 50% debt financed. If the firm's corporate tax rate is 0.2, the typical personal tax rate of an investor in the firm's stock is 0.5, and the typical tax rate for an investor in the firm's debt is 0.6. What will be the new value of the firm under the MM THEORY WITH CORPORATE TAXES BUT NO POSSIBILITY OF BANKRUPTCY.
Answer #1

Value of an all equity financed company, V is given by:

V = (1 - Personal Tax on Equity)*(1 - Corporate tax)*Value of unlevered firm

V = (1 - 0.5)*(1 - 0.2)*425100 =0.5*0.8*425100

=$170040

Now, Value of Levered firm, Vd

Vd = {V}+{V}*{P}*(1 - ( 1-{Tc})*(1-{Ts})/(1-{Td}) )/100

P is proportion of debt = 0.50

Tc is corporate tax

Ts is tax on stock market for investor

Td is tax on debt for investor

Vd = 170040+ (170040 * 0.50 * (1 - (1 - 0.2) * (1 - 0.5)/(1 - 0.6))

Vd = 170040 + 170040*0.50* (1-(0.8*0.5)/0.4

= 170040+ 170040*0.5(1-1)

Vd = 170040 + (0)= 170040